Trace the curve x = (y – 1) (y – 2) (y – 5)


1
Expert's answer
2021-07-09T08:47:03-0400

1. Equation of the curve : 

"x=(y-1)(y-2)(y-5)"




"x=(y-1)(y^2-7y+10)"

"x=y^3-8y^2+17y-10"

Domain: "(-\infin, \infin)."

Range: "(-\infin, \infin)."


2. Symmetry

"-y:"


"x=(-y-1)(-y-2)(-y-5)"

"x=-(y+1)(y+2)(y+5)"

The curve is not symmetrical about "x" -axis.


"-x:"


"-x=(y-1)(y-2)(y-5)"

"x=-(y-1)(y-2)(y-5)"

The curve is not symmetrical about "y" -axis.


"-x, -y" :



"-x=(-y-1)(-y-2)(-y-5)"

"x=(y+1)(y+2)(y+5)"

The curve is not symmetrical about the origin.


Interchange "x" and "y:"


"y=(x-1)(x-2)(x-5)"

The curve is not symmetrical about the line "y=x."


The curve is not symmetric with respect to any line or axis.


3. Intersection(s)

"Oy: x=0"


"0=(y-1)(y-2)(y-5)"

"Point(0,1), Point(0, 2), Point(0, 5)"



"Ox: y=0"

"x=(0-1)(0-2)(0-5)"

"Point(-10,0)."


4. Tangent at Origin :

The curve does not go through the origin. So there is no tangent at the origin.


5. There does not exist any horizontal or vertical asymptote.

"y\to-\infin" as "x\to-\infin"


"y\to\infin" as "x\to\infin"


There is no slant (oblique) asymptote.


6. Local maximum or minimum.


"x=y^3-8y^2+17y-10"

Differentiate both sides with respect to "y"

"x_y'=3y^2 -16y+17"

Critical number(s)

"x'=0=>3y^2 -16y+17=0"

"y=\dfrac{8\pm\sqrt{13}}{3}"

"Point(-6.065, 3.869)" is a local minimum.


"Point(0.879, 1.465)" is a local maximum.


7. Sketch the graph